# Equivariant Machine Learning on Graphs with Nonlinear Spectral Filters

### v1

@@ B.1 Semi-Supervised Node Classification
Datasets. We consider six datasets in node classification tasks, including Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Actor. The detailed statistics of the node classification benchmarks are summarized in Tab. LABEL:tab:node_classification_statistics. For datasets splitting on citation graphs (Cora, Citeseer, and Pubmed), we follow the standard splits from [Yang et al., 2016], using 20 nodes per class for training, 500 nodes for validation, and 1000 nodes for testing. For heterophilic graphs (Chameleon, Squirrel, and Actor), we adopt the sparse splitting method from [Chien et al., 2020, He et al., 2022], allocating 2.5% of samples for training, 2.5% for validation, and 95% for testing. The classification quality is assessed by computing the average classification accuracy across 10 random splits, along with a 95% confidence interval.
Baselines. For GCN, GAT, SAGE, ChebNet, ARMA, and APPNP, we use the implementations from the PyTorch Geometric library [Fey and Lenssen, 2019]. For other baselines, we use the implementations released by the respective authors.
Hyperparameters Settings. The hidden dimension is set to be either 64 or 128 for all models and datasets. We implement our proposed Node-level NLSFs using PyTorch and optimize the model with the Adam optimizer [Kingma and Ba, 2014]. To determine the optimal dropout probability, we search within the range $[0,0.9]$ in increments of $0.1$. The learning rate is examined within the set $\{1e^{-1},5e^{-2},1e^{-2},5e^{-3},1e^{-3}\}$. We explore weight decay values within the set $\{1e^{-2},5e^{-3},1e^{-3},5e^{-4},1e^{-4},5e^{-5},1e^{-5},5e^{-6},1e^{-6},0.0\}$. Furthermore, the number of layers is varied from 1 to 10. The number of leading eigenvectors $J$ in Index NLSFs is set within $[1,1e^{2}]$. The decay rate in the Value NLSFs is determined using dyadic sampling within the set $\{\frac{1}{4},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{3}{4}\}$, the sampling resolution $S$ within $[1,1e^{2}]$, and the number of the bands in Value NLSFs $K\leq S$. For hyperparameter optimization, we conduct a grid search using Optuna [Akiba et al., 2019] for each dataset. An early stopping criterion is employed during training, stopping the process if the validation loss does not decrease for 200 consecutive epochs.
@@
B.2 Graph Classification
@@
Datasets. We consider eight datasets [Morris et al., 2020] for graph classification tasks, including five bioinformatics: MUTAG, PTC, NCI1, ENZYMES, and PROTEINS, and three social network datasets: IMDB-B, IMDB-M, and COLLAB. The detailed statistics of the graph classification benchmarks are summarized in Tab. LABEL:tab:graph_classification_statistics. We use the random split from [Velickovic et al., 2019, Ying et al., 2018, Ma et al., 2019, Zhu et al., 2021b], using 80% for training, 10% for validation, and 10% for testing. This process is repeated 10 times, and we report the average performance and standard deviation.
Baselines. For GCN, GAT, SAGE, ChebNet, ARMA, APPNP, and DiffPool, we use the implementations from the PyTorch Geometric library [Fey and Lenssen, 2019]. For other baselines, we use the implementations released by the respective authors.
Hyperparameters Settings. The dimension of node representations is set to 128 for all methods and datasets. We implement the proposed Pooling-NLSFs and Graph-level NLSFs using PyTorch and optimize the model with the Adam optimizer [Kingma and Ba, 2014]. A readout function is applied to aggregate the node representations for each graph, utilizing mean, add, max, or RMS poolings. The learning rate and weight decay are searched within $\{1e^{-1},1e^{-2},1e^{-3},1e^{-4},1e^{-5}\}$, the pooling ratio within $[0.1,0.9]$ with step $0.1$, the number of layers within $[1,10]$ with step $1$, the number of leading eigenvectors $J$ in Index NLSFs within $[1,1e^{2}]$, the decay rate in the Value NLSFs using dyadic sampling within $\{\frac{1}{4},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{3}{4}\}$, the sampling resolution $S$ within $[1,1e^{2}]$, and the number of the bands in Value NLSFs within $K\leq S$. The graph-level representation is then fed into an MLP with a (log)softmax classifier, using a cross-entropy loss function for predictions over the labels. Specifically, the MLP consists of three fully connected layers with 256, 128, and 64 neurons, respectively, followed by a (log)softmax classifier. We conduct a grid search on the hyperparameters for each dataset using Optuna [Akiba et al., 2019]. An early stopping criterion is employed during training, stopping the process if the validation loss does not decrease for 100 consecutive epochs.
Appendix C Additional Experimental Results
@@
C.1 Semi-Supervised Node Classification Following [He et al., 2022] Protocol
Table 6: Semi-supervised node classification accuracy with random split, following the experimental protocol established by [He et al., 2022]. Results marked with are taken from [He et al., 2022].
@@
We present additional experimental results for semi-supervised node classification using random splits, adhering to the protocol established by [He et al., 2022]. The results, summarized in Tab. LABEL:tab:node_classification_accuracy_following_chebnetII_setup, demonstrate the classification accuracy across six benchmark datasets: Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Actor. Our $\mathtt{att}$-Node-level NLSFs achieve the highest accuracy on four out of the six datasets, outperforming other models significantly. On Pubmed, it records a close second-highest accuracy, slightly behind APPNP. Our method achieves a competitive second place in the Actor dataset. $\mathtt{att}$-Node-level NLSFs demonstrate substantial improvements, particularly in challenging datasets like Chameleon and Squirrel. The comparison models, including GCN, GAT, SAGE, ChebNet, ChebNetII, CayleyNet, APPNP, GPRGNN, and ARMA, varied in their effectiveness, with ChebNetII and APPNP also showing strong results on several datasets. These findings highlight the efficacy of the $\mathtt{att}$-Node-level NLSFs in semi-supervised node classification tasks.
C.2 Node Classification on Filtered Chameleon and Squirrel Datasets in Dense Split Setting
Table 7: Full Performance comparison of node classification on original and filtered Chameleon and Squirrel datasets in dense split setting. The baseline results used for comparison are taken from [Platonov et al., 2023].
@@ C.2 Node Classification on Filtered Chameleon and Squirrel Datasets in Dense Split Setting
Following [Platonov et al., 2023], we conduct additional experiments on the original and filtered Chameleon and Squirrel datasets in the dense split setting. We use the same random splits as in [Platonov et al., 2023], dividing the datasets into 48% for training, 32% for validation, and 20% for testing. We compare the Node-level NLSFs using Laplacian attention with GCN [Kipf and Welling, 2016], SAGE [Hamilton et al., 2017], GAT [Veličković et al., 2017], GT [Shi et al., 2020], H2GCN [Zhu et al., 2020], CPGNN [Zhu et al., 2021a], GPRGNN [Chien et al., 2020], FSGNN [Maurya et al., 2022], GloGNN [Li et al., 2022], FAGCN [Bo et al., 2021], GBKGNN [Du et al., 2022], JacobiConv [Wang and Zhang, 2022], and ResNet [He et al., 2016] with GNN models [Platonov et al., 2023]. The study by [Zhu et al., 2020] demonstrates the benefits of separating ego- and neighbor-embeddings in the GNN aggregation step when dealing with heterophily. Therefore, [Platonov et al., 2023] also adopts this approach for the GNN aggregation step in GAT and GT models, denoted as “sep.” The baseline results used for comparison are taken from [Platonov et al., 2023]. Tab. 7 presents the full performance comparison on the original and filtered datasets. Note that Tab. 7 is the same as Tab. 2 but with more baseline methods. $\mathtt{att}$-Node-level NLSFs achieve the highest accuracy on both the filtered Chameleon and Squirrel datasets. Additionally, $\mathtt{att}$-Node-level NLSFs demonstrate strong performance on the original Chameleon dataset, achieving the highest accuracy and the second-highest accuracy on the original Squirrel dataset. $\mathtt{att}$-Node-level NLSFs show less sensitivity to node duplicates and exhibit stronger generalization ability, further validating the reliability of the Chameleon and Squirrel datasets in the dense split setting.
C.3 Ablation Study on $\mathtt{att}$-Node-level NLSFs, $\mathtt{att}$-Graph-NLSF, and $\mathtt{att}$-Pooling-NLSFs
@@
The ablation study of $\mathtt{att}$-Node-level NLSPs for node classification is summarized in Tab. LABEL:tab:node_classification_nlsfp_ablation. We investigate the performance on six node classification benchmarks as in Sec. 5, including Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Actor. Tab. LABEL:tab:node_classification_nlsfp_ablation shows that using the graph Laplacian $\mathbf{L}$ in Index NLSFs (denoted as $\Theta_{\text{ind}}(\mathbf{L},\cdot)$) and the normalized graph Laplacian $\mathbf{N}$ in Value NLSFs (denoted as $\Theta_{\text{val}}(\mathbf{N},\cdot)$ ) has superior node classification accuracy compared to using the normalized graph Laplacian $\mathbf{N}$ in Index NLSFs (denoted as $\Theta_{\text{ind}}(\mathbf{N},\cdot)$ ) and the graph Laplacian $\mathbf{L}$ in Value NLSFs (denoted as $\Theta_{\text{val}}(\mathbf{L},\cdot)$). This is in line with our theoretical findings in App. A.5. Moreover, the $\mathtt{att}$-Node-level NLSFs using the Laplacian attention $\mathtt{att}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ yield the highest accuracies across all datasets, corroborating the findings in Sec. 3.5. We also note that without Laplacian attention, the Node-level NLSFs ($\Theta_{\text{ind}}(\mathbf{L},\cdot)$ and $\Theta_{\text{val}}(\mathbf{N},\cdot)$) alone still achieve more effective classification performance compared to existing baselines, as shown in Tab. LABEL:tab:node_classification_accuracy.
@@
The ablation study of $\mathtt{att}$-Pooling-NLSFs for graph classification is reported in Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling. Similar to Tab. LABEL:tab:graph_classification_nlsfp_ablation_graph_nlsf, we consider eight graph classification benchmarks: MUTAG, PTC, ENZYMES, PROTEINS, NCI1, IMDB-B, IMDB-M, and COLLAB. Tabl. LABEL:tab:graph_classification_nlsfp_ablation_pooling demonstrates that using the graph Laplacian $\mathbf{L}$ in Index Pooling-NLSFs (denoted as $\Theta_{\text{ind}}^{\mathtt{P}}(\mathbf{L},\cdot)$) and the normalized graph Laplacian $\mathbf{N}$ in Value Pooling-NLSFs (denoted as $\Theta_{\text{val}}^{\mathtt{P}}(\mathbf{N},\cdot)$) achieves superior graph classification accuracy compared to using the normalized graph Laplacian $\mathbf{N}$ in Index Pooling-NLSFs (denoted as $\Theta_{\text{ind}}^{\mathtt{P}}(\mathbf{N},\cdot)$) and the graph Laplacian $\mathbf{L}$ in Value Pooling-NLSFs (denoted as $\Theta_{\text{val}}^{\mathtt{P}}(\mathbf{L},\cdot)$). This finding aligns with our theoretical results in App. A.5. Moreover, the $\mathtt{att}$-Pooling-NLSFs using the Laplacian attention $\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ yield the highest accuracies across all datasets, corroborating the findings in Sec. 3.5. In addition, $\mathtt{att}$-Pooling-NLSFs consistently outperform $\mathtt{att}$-Graph-level NLSFs as shown in Tab. LABEL:tab:graph_classification_nlsfp_ablation_graph_nlsf, indicating that the node features learned in our Node-level NLSFs representation are more expressive. This supports our theoretical findings in Sec. 4.3.
@@ C.4 Runtime Analysis
C.5 Diagonal Nonlinear Spectral Filter

### v2

@@ E.1 Semi-Supervised Node Classification
Datasets. We consider six datasets in node classification tasks, including Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Actor. The detailed statistics of the node classification benchmarks are summarized in Tab. LABEL:tab:node_classification_statistics. For datasets splitting on citation graphs (Cora, Citeseer, and Pubmed), we follow the standard splits from [100], using 20 nodes per class for training, 500 nodes for validation, and 1000 nodes for testing. For heterophilic graphs (Chameleon, Squirrel, and Actor), we adopt the sparse splitting method from [16, 41], allocating 2.5% of samples for training, 2.5% for validation, and 95% for testing. The classification quality is assessed by computing the average classification accuracy across 10 random splits, along with a 95% confidence interval.
Baselines. For GCN, GAT, SAGE, ChebNet, ARMA, and APPNP, we use the implementations from the PyTorch Geometric library [28]. For other baselines, we use the implementations released by the respective authors.
Hyperparameters Settings. The hidden dimension is set to be either 64 or 128 for all models and datasets. We implement our proposed Node-level NLSFs using PyTorch and optimize the model with the Adam optimizer [48]. To determine the optimal dropout probability, we search within the range $[0,0.9]$ in increments of $0.1$. The learning rate is examined within the set $\{1e^{-1},5e^{-2},1e^{-2},5e^{-3},1e^{-3}\}$. We explore weight decay values within the set $\{1e^{-2},5e^{-3},1e^{-3},5e^{-4},1e^{-4},5e^{-5},1e^{-5},5e^{-6},1e^{-6},0.0\}$. Furthermore, the number of layers is varied from 1 to 10. The number of leading eigenvectors $J$ in Index NLSFs is set within $[1,1e^{2}]$. The decay rate in the Value NLSFs is determined using dyadic sampling within the set $\{\frac{1}{4},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{3}{4}\}$, the sampling resolution $S$ within $[1,1e^{2}]$, and the number of the bands in Value NLSFs $K\leq S$. For hyperparameter optimization, we conduct a grid search using Optuna [1] for each dataset. An early stopping criterion is employed during training, stopping the process if the validation loss does not decrease for 200 consecutive epochs.
E.2 Graph Classification
@@
Datasets. We consider eight datasets [72] for graph classification tasks, including five bioinformatics: MUTAG, PTC, NCI1, ENZYMES, and PROTEINS, and three social network datasets: IMDB-B, IMDB-M, and COLLAB. The detailed statistics of the graph classification benchmarks are summarized in Tab. LABEL:tab:graph_classification_statistics. We use the random split from [94, 101, 67, 104], using 80% for training, 10% for validation, and 10% for testing. This process is repeated 10 times, and we report the average performance and standard deviation.
Baselines. For GCN, GAT, SAGE, ChebNet, ARMA, APPNP, and DiffPool, we use the implementations from the PyTorch Geometric library [28]. For other baselines, we use the implementations released by the respective authors.
Hyperparameters Settings. The dimension of node representations is set to 128 for all methods and datasets. We implement the proposed Pooling-NLSFs and Graph-level NLSFs using PyTorch and optimize the model with the Adam optimizer [48]. A readout function is applied to aggregate the node representations for each graph, utilizing mean, add, max, or RMS poolings. The learning rate and weight decay are searched within $\{1e^{-1},1e^{-2},1e^{-3},1e^{-4},1e^{-5}\}$, the pooling ratio within $[0.1,0.9]$ with step $0.1$, the number of layers within $[1,10]$ with step $1$, the number of leading eigenvectors $J$ in Index NLSFs within $[1,1e^{2}]$, the decay rate in the Value NLSFs using dyadic sampling within $\{\frac{1}{4},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{3}{4}\}$, the sampling resolution $S$ within $[1,1e^{2}]$, and the number of the bands in Value NLSFs within $K\leq S$. The graph-level representation is then fed into an MLP with a (log)softmax classifier, using a cross-entropy loss function for predictions over the labels. Specifically, the MLP consists of three fully connected layers with 256, 128, and 64 neurons, respectively, followed by a (log)softmax classifier. We conduct a grid search on the hyperparameters for each dataset using Optuna [1]. An early stopping criterion is employed during training, stopping the process if the validation loss does not decrease for 100 consecutive epochs.
Appendix F Additional Experimental Results
@@
F.1 Semi-Supervised Node Classification Following [41] Protocol
We present additional experimental results for semi-supervised node classification using random splits, adhering to the protocol established by [41]. The results, summarized in Tab. LABEL:tab:node_classification_accuracy_following_chebnetII_setup, demonstrate the classification accuracy across six benchmark datasets: Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Actor. Our $\mathtt{att}$-Node-level NLSFs achieve the highest accuracy on four out of the six datasets, outperforming other models significantly. On Pubmed, it records a close second-highest accuracy, slightly behind APPNP. Our method achieves a competitive second place in the Actor dataset. $\mathtt{att}$-Node-level NLSFs demonstrate substantial improvements, particularly in challenging datasets like Chameleon and Squirrel. The comparison models, including GCN, GAT, SAGE, ChebNet, ChebNetII, CayleyNet, APPNP, GPRGNN, and ARMA, varied in their effectiveness, with ChebNetII and APPNP also showing strong results on several datasets. These findings highlight the efficacy of the $\mathtt{att}$-Node-level NLSFs in semi-supervised node classification tasks.
F.2 Node Classification on Filtered Chameleon and Squirrel Datasets in Dense Split Setting
Table 7: Semi-supervised node classification accuracy with random split, following the experimental protocol established by [41]. Results marked with are taken from [41].
@@ F.2 Node Classification on Filtered Chameleon and Squirrel Datasets in Dense Split Setting
Table 8: Full Performance comparison of node classification on original and filtered Chameleon and Squirrel datasets in dense split setting. The baseline results used for comparison are taken from [77].
@@
Following [77], we conduct additional experiments on the original and filtered Chameleon and Squirrel datasets in the dense split setting. We use the same random splits as in [77], dividing the datasets into 48% for training, 32% for validation, and 20% for testing. We compare the Node-level NLSFs using Laplacian attention with GCN [49], SAGE [39], GAT [93], GT [88], H2GCN [103], CPGNN [102], GPRGNN [16], FSGNN [69], GloGNN [62], FAGCN [8], GBKGNN [23], JacobiConv [96], and ResNet [40] with GNN models [77]. The study by [103] demonstrates the benefits of separating ego- and neighbor-embeddings in the GNN aggregation step when dealing with heterophily. Therefore, [77] also adopts this approach for the GNN aggregation step in GAT and GT models, denoted as “sep.” The baseline results used for comparison are taken from [77]. Tab. 8 presents the full performance comparison on the original and filtered datasets. Note that Tab. 8 is the same as Tab. 2 but with more baseline methods. $\mathtt{att}$-Node-level NLSFs achieve the highest accuracy on both the filtered Chameleon and Squirrel datasets. Additionally, $\mathtt{att}$-Node-level NLSFs demonstrate strong performance on the original Chameleon dataset, achieving the highest accuracy and the second-highest accuracy on the original Squirrel dataset. $\mathtt{att}$-Node-level NLSFs show less sensitivity to node duplicates and exhibit stronger generalization ability, further validating the reliability of the Chameleon and Squirrel datasets in the dense split setting.
F.3 Ablation Study on $\mathtt{att}$-Node-level NLSFs, $\mathtt{att}$-Graph-NLSF, and $\mathtt{att}$-Pooling-NLSFs
@@
The ablation study of $\mathtt{att}$-Node-level NLSPs for node classification is summarized in Tab. LABEL:tab:node_classification_nlsfp_ablation. We investigate the performance on six node classification benchmarks as in Sec. 5, including Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Actor. Tab. LABEL:tab:node_classification_nlsfp_ablation shows that using the graph Laplacian $\mathbf{L}$ in Index NLSFs (denoted as $\Theta_{\text{ind}}(\mathbf{L},\cdot)$) and the normalized graph Laplacian $\mathbf{N}$ in Value NLSFs (denoted as $\Theta_{\text{val}}(\mathbf{N},\cdot)$ ) has superior node classification accuracy compared to using the normalized graph Laplacian $\mathbf{N}$ in Index NLSFs (denoted as $\Theta_{\text{ind}}(\mathbf{N},\cdot)$ ) and the graph Laplacian $\mathbf{L}$ in Value NLSFs (denoted as $\Theta_{\text{val}}(\mathbf{L},\cdot)$). This is in line with our theoretical findings in App. D.5. Moreover, the $\mathtt{att}$-Node-level NLSFs using the Laplacian attention $\mathtt{att}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ yield the highest accuracies across all datasets, corroborating the findings in Sec. 3.5. We also note that without Laplacian attention, the Node-level NLSFs ($\Theta_{\text{ind}}(\mathbf{L},\cdot)$ and $\Theta_{\text{val}}(\mathbf{N},\cdot)$) alone still achieve more effective classification performance compared to existing baselines, as shown in Tab. LABEL:tab:node_classification_accuracy.
@@
The ablation study of $\mathtt{att}$-Pooling-NLSFs for graph classification is reported in Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling. Similar to Tab. LABEL:tab:graph_classification_nlsfp_ablation_graph_nlsf, we consider eight graph classification benchmarks: MUTAG, PTC, ENZYMES, PROTEINS, NCI1, IMDB-B, IMDB-M, and COLLAB. Tabl. LABEL:tab:graph_classification_nlsfp_ablation_pooling demonstrates that using the graph Laplacian $\mathbf{L}$ in Index Pooling-NLSFs (denoted as $\Theta_{\text{ind}}^{\mathtt{P}}(\mathbf{L},\cdot)$) and the normalized graph Laplacian $\mathbf{N}$ in Value Pooling-NLSFs (denoted as $\Theta_{\text{val}}^{\mathtt{P}}(\mathbf{N},\cdot)$) achieves superior graph classification accuracy compared to using the normalized graph Laplacian $\mathbf{N}$ in Index Pooling-NLSFs (denoted as $\Theta_{\text{ind}}^{\mathtt{P}}(\mathbf{N},\cdot)$) and the graph Laplacian $\mathbf{L}$ in Value Pooling-NLSFs (denoted as $\Theta_{\text{val}}^{\mathtt{P}}(\mathbf{L},\cdot)$). This finding aligns with our theoretical results in App. D.5. Moreover, the $\mathtt{att}$-Pooling-NLSFs using the Laplacian attention $\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ yield the highest accuracies across all datasets, corroborating the findings in Sec. 3.5. In addition, $\mathtt{att}$-Pooling-NLSFs consistently outperform $\mathtt{att}$-Graph-level NLSFs as shown in Tab. LABEL:tab:graph_classification_nlsfp_ablation_graph_nlsf, indicating that the node features learned in our Node-level NLSFs representation are more expressive. This supports our theoretical findings in Sec. 4.3.
@@ F.4 Runtime Analysis
F.5 Scalability to Large-Scale Datasets
Table 13: Experimental results on large heterophilic graphs. The results for BernNet, ChebNet, ChebNetII, and GPRGNN are taken from [41], while the results for OptBasisGNN are taken from [38]. All other competing results are taken from [63].
Penn94 Pokec Genius Twitch-Gamers Wiki
GCN 82.47$\pm$0.3 75.45$\pm$0.2 87.42$\pm$0.4 62.18$\pm$0.3 OOM
LINK 80.79$\pm$0.5 80.54$\pm$0.0 73.56$\pm$0.1 64.85$\pm$0.2 57.11$\pm$0.3
LINKX 84.71$\pm$0.5 82.04$\pm$0.1 90.77$\pm$0.3 66.06$\pm$0.2 59.80$\pm$0.4
GPRGNN 83.54$\pm$0.3 80.74$\pm$0.2 90.15$\pm$0.3 62.59$\pm$0.4 58.73$\pm$0.3
ChebNet 82.59$\pm$0.3 72.71$\pm$0.7 89.36$\pm$0.3 62.31$\pm$0.4 OOM
ChebNetII 84.86$\pm$0.3 82.33$\pm$0.3 90.85$\pm$0.3 65.03$\pm$0.3 60.95$\pm$0.4
BernNet 83.26$\pm$0.3 81.67$\pm$0.2 90.47$\pm$0.3 64.27$\pm$0.3 59.02$\pm$0.3
OptBasisGNN 84.85$\pm$0.4 82.83$\pm$0.0 90.83$\pm$0.1 65.17$\pm$0.2 61.85$\pm$0.0
$\mathtt{att}$-Node-level NLSFs 85.19$\pm$0.3 82.96$\pm$0.1 91.24$\pm$0.1 65.97$\pm$0.2 62.44$\pm$0.3
To demonstrate the scalability of our method, we conduct additional tests on five large heterophilic graphs: Penn94, Pokec, Genius, Twitch-Gamers, and Wiki datasets from [63]. The experimental setup is in line with previous work by [16, 63, 41]. We use the same hyperparameters for our NLSFs as reported in App. E. Tab. LABEL:tab:large_dataset presents the classification accuracy. We see that NLSFs outperform the competing methods on the Penn94, Pokec, Genius, and Wiki datasets. For the Twitch-Gamers dataset, NLSFs yield the second-best results. Our additional experiments show that our method could indeed scale to handle large-scale graphs effectively.
F.6 Index-by-Index Index-NLSFs and Band-by-Band Value-NLSFs